OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!!!!!Find the product. -1/2times(-8)

OpenStudy (anonymous):

ok ill try to help

OpenStudy (jackky):

4

OpenStudy (anonymous):

-.-

OpenStudy (jackky):

im pretty sure

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

A. –16 B. –4 C. 4 D. 16

OpenStudy (causticsyndicalist):

Last I checked, direct answers were against the Code of Conduct. (~._.)~

OpenStudy (jackky):

c

OpenStudy (anonymous):

tnx

OpenStudy (jackky):

ya welcome

OpenStudy (anonymous):

Ya it's 4 =-1 /2×8/1 =(-1×-8)/(2×1) =8/2 =4/1 =4